Diffusion Morning Report 07/25/06

Ted Berg tberg at stanford.edu
Tue Jul 25 06:45:59 PDT 2006


Good Morning All,
   Loooking good this morning. Nitride temp looks ok now will keep it 
yellow to observe. Gas testing for new Epi went well. GeH4 tie in
is almost completed. Burn boxwas a bit high this morning-cleaning. AMAT 
Epi install moving along.  Tylan 2 has a door close issue-checking it out.
Everything else looks OK for now.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	7/25/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	yellow 	loader issue
	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	1.8/30 	  	 
Tylan 8-SiGe 	down 	GEH4 tie in, tube pull completed 	0/30 	  	 
Tylan 9-Poly 	Up 	  	17.91/30 	  	 
Tylan 10-Nitride 	down 	checking 	19.99/30 	  	 
Tylan 11-BPSG-LTO 	Up 	  	14.2/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
  	  	  	  	  	 
Epi 	Up 	  	  	  	 
  	  	  	  	  	 
Burn Box 	Up 	 cleaning this morning
	  	  	 
ZYGO 	up 	  	   	  	 
  	  	  	  	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
  	  	  	  	  	 
SiH4 for furnaces 	  	A on line 100 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060725/8bccc498/attachment.html>


More information about the diffusion mailing list